OpenStudy (anonymous):

What is the graph of the absolute value equation? y = –1/2|x|

OpenStudy (anonymous):

Make a list of x values, substitute them into your equation and your answers is the y value. absolute values around the x just mean that even a negative number will be positive. ex: x=1 y=-1/2(1)

OpenStudy (anonymous):

And then use your points to graph it.

OpenStudy (anonymous):

@completeidiot

OpenStudy (anonymous):

yea to understand what multiples do to the absolute value graph, i would suggest what ottomaticc is saying just plot 2 instances 1/2|x| and 2|x|

OpenStudy (anonymous):

but lets dissect this anyways lets start by looking at the negative sign what does the negative sign do the graph?

OpenStudy (anonymous):

this one I think is different from the other ones we did

OpenStudy (anonymous):

makes it negative so the lines would go downward wait let me attacth pics

OpenStudy (anonymous):

ok based off the information given from the negative sign, which 2 graphs are correct?

OpenStudy (anonymous):

one and two

OpenStudy (anonymous):

right, so now we just look at the 1/2 any idea what that will do to the graph? a good way to think about it is by thinking about the slope of a line

OpenStudy (anonymous):

no I dn't

OpenStudy (anonymous):

do you remember slope intercept form? y=mx+b

OpenStudy (anonymous):

eh not so much

OpenStudy (anonymous):

Rise over run, up one over two

OpenStudy (anonymous):

m is the slope 

OpenStudy (anonymous):

in this case, your slope would be 1/2, think of |x| as just x for now which of the 2 graph looks like the y="1/2 x"

OpenStudy (anonymous):

capture 2

OpenStudy (anonymous):

??

OpenStudy (anonymous):

no , meant my graphs, not yours

OpenStudy (anonymous):

oh then I don't know

OpenStudy (anonymous):

m= 1/2 since 1/2 is not greater than 1, then its no the first graph that means the slope would look like my 2nd graph

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now you know what the slope vaguely looks like, compare that to the 2 graphs you have

OpenStudy (anonymous):

capture 1 I think it looks more similar :)

OpenStudy (anonymous):

but wat do I know I hate math my brain can't keep up")

OpenStudy (anonymous):

right now lets check our answer lets pick a random point on the graph other than (0,0), say (6,-3) and plug the values into y = –1/2|x| -3 = –1/2|6| -3=-1/2*6 -3=-3 so you are right

OpenStudy (anonymous):

oh alright thank you sooo much!!

OpenStudy (anonymous):

so now you know when you have a multiple infront of the absolute value that is between 0 and 1, the graph expands and when the multiple in front of the absolute value is greater than 1, the graph shrinks

OpenStudy (anonymous):

yup! thanks!!!!